# Ligament mediated aerobreakup of a droplet

### v1

@@ III.3 Ligament-mediated droplet generation
As explained earlier, the sub-secondary breakup processes ends as the undulation disintegrates into daughter droplets. However, the termination entails a final ligament-mediated mechanism. The corrugated ligaments portrays an inertio-capillary system with liquid inertia and surface tension forces dominating the dynamics. The ligament can be considered as a collection of a protoblobs where the evolution ensure two possibilities: either the blobs separate through RPI mechanism or they coalesce into a bigger blob. Recent studies [8, 10, 27] considered these effects and emphasized the role of corrugations on this delicate balance between breakup and coalescence. Hence a correlation between the ligament shape and the final daughter droplet size distribution has been established in the form of a Gamma distribution [8] or possibly a compound gamma distribution [9]
@@
where $\Gamma$ is the gamma function, $\mathcal{K}$ is the modified Bessel function of the second kind and the coefficients $m$ and $n$ represent ligament size distribution and corrugations. $n=\langle d_{c}\rangle^{2}/(\langle d_{c}^{2}\rangle-\langle d_{c}\rangle^{2})$ is deduced by decomposing a corrugated ligament into blobs matching local ligament diameter $d_{c}$ as illustrated in Fig. 7. $m=\langle l\rangle^{2}/(\langle l^{2}\rangle-\langle l\rangle^{2})$ is deduced by taking the mean diameters $l=\langle d_{c}\rangle$ for various ligaments. Considering a sample of ligaments spread across space and time with different imposed gas flow We, this corrugation factor was determined mostly to be in the range spanning $n=3-8$. The values close to $n=4$ were identified as the most frequently observed across various $We$ flows. This corresponds to extremely corrugated ligaments and $n=4$ is a practical limit on the coefficient considering various physical constraints [10]. Also a very low value of $m=3$ (corresponding to the fit) represent a broad ligament size ($l$) distribution, as observed from the experiments as well. The invariance with respect to $We$ number supports the saturation of this coefficient to the theoretical limit. Although the sub-secondary breakup process involves the ligament production through various mechanism, the ligaments collectively are postulated to exhibit extreme levels of corrugations, primarily due to the highly potent disruptive aerodynamic forces at extreme We. The ligaments are expected to scale down to smaller average diameters as $We$ is increased (as observed in the experiments in Fig. 5), but the shape factor is already saturated to the limit and depicts values near $n=4$. To elucidate this hypothesis, we will consider the estimated size distribution in the subsequent sections.
@@ III.4 Droplet size distribution
The unsteady imposition arising from the open nozzle-blast wave induced flows [28], the dynamics are expected to be transient. However the droplet aerobreakup intrinsically is a transient phenomenon. The recurrent wave cascade continuously strips away the liquid from the parent droplet mass and generates droplets through sub-secondary breakup processes. Hence the droplets are produced with different characteristics over a period of time. To capture this effect, droplet size measurements were performed considering the time stamp ($t^{*}$) of their appearance in the measurement zone. This zone was located $30mm$ downstream from the droplet for better visibility [14].
@@
Figure 8: (a) Overall droplet size distribution PDFs at different We. (b) Temporal evolution of PDFs represented a contour plot, where the breakup event is sub divided into 5 equal periods. (c) Temporal evolution of Normalised PDFs (using average droplet diameter $\langle d\rangle$), depicting self-similar behaviour in time. Here $\tau_{b}$=centre of each time period/total breakup time.
@@
All the sampled droplets, when consolidated, gave the overall size distributions for different $We$ flow. However, the appearance of the first and the last droplet in the zone reflects the breakup time period $t_{b}^{*}$, which can be subdivided into equal sub-periods. The assessment of size distributions in these periods enables the determination of transient characteristics of the system. The time scale $t^{*}$, measured from the appearance of the first droplet, is then normalized using the breakup period $t_{b}^{*}$ such that $\tau_{b}=t^{*}/t_{b}^{*}$. The overall distributions as probability density functions (PDFs) are illustrated in Fig. 8a. The average droplet size decreases as $We$ is increased. The transient PDFs are represented as a contour plot depicting a transition to observing larger droplet sizes as time progresses in Fig. 8b. However, if we normalize the PDFs with corresponding average droplet size $d_{10}^{*}$, we observe the different time periods follows same distributions. The same effects are visible in all $We$ discussed in this study as illustrated in Fig. 8c.
@@
For more reliable statistics with a larger sample size in each period, the breakup time is split into 3 parts, and the corresponding normalized PDFs along with global distributions are presented in in Fig. 9a along with the compound gamma and log-normal distribution fits. They almost overlap, especially till $d/\langle d\rangle\approx 6$. The compound gamma distribution aligns with the earlier prediction $n=4$, deduced from the ligament-mediated mechanism. The deviation is observed consistently at the tails, which might be due to the effects like (i) The unsteady decaying airflow with lower effective $We$ at later stages generating larger droplets produced from the residual parent droplet and (ii) coalescence of daughter droplets. To have a better understanding, we require unsteady destabilization models and better measurement capabilities.
@@
Since the normalized PDFs follow a common underlying distribution, the prediction of average diameter is ideally sufficient to determine actual PDFs. A previous study [29] assumed a chaotic isotropic deformation of the liquid mass to analytically deduce size distributions in explosive breakups. We extend the same to present catastrophic aerobreakup. The imposed disruptive energy $E$ induces a deformation rate of $\dot{\gamma}$ to the droplet of initial radius $r_{0}$, such that $(1/2)\rho_{l}r_{0}^{2}(\dot{\gamma}r_{0})^{2}\sim E$. Assuming chaotic random dispersal of energy [29], the same deformation rate is extended to individual daughter droplets with characteristic size $r_{f}$ (radius), such that the associated incremental surface energy scales as $(1/2)\rho_{l}r_{f}^{2}(\dot{\gamma}r_{0})^{2}\sim\sigma r_{f}^{2}$. Simplifying these equations where $E\sim\rho_{a}u_{a}^{2}r_{0}^{3}$ for aerobreakup, we get $(r_{f}/r_{0})\sim{We}^{-1/3}$. Suspecting $d_{10}\sim d_{f}$, we have
@@
This scaling relationship is validated in Fig. 9b. Although the observed breakup is not completely random, the hypothesis holds. Deviations from this are expected at lower We, where the breakup is evidently regular in geometry and anisotropic. A similar scaling law for average droplet diameters was deduced for flat-fan sprays, using the unstable wave-mediated sheet disintegration mechanism [9].
A temporal evolution of average droplet diameters $d_{10}^{*}$ are also presented in Fig. 10a, where global averages $d_{10}$ is used to normalization, depicting a similar trend for different We, where smaller droplets appear first (also see Fig. 8b). The possible mechanisms include: (i) unsteady impulsive airflow with higher effective $We$ generating smaller droplets during early interaction (ii) size dependent migration where smaller droplets with lower inertia reach the measurement zone first. The number of daughter droplets $N$ can be also deduced extending earlier analysis
@@
We consider the number of droplets produced per unit measurement volume i.e. the number concentration $n=N/V_{meas}$ for unbiased estimates of droplet count [14, 30]. Temporal evolution of the normalized version $n/We$ is plotted in Fig. 10b illustrating the validity of aforementioned scales. A similar trend is observed for different We, where a huge number of droplets are generated in the early stages. Culminating both the ideas (temporal evolution of $d_{10}^{*}$ and $n$), we deduce that the droplets with smaller droplets appear first in large quantities. Since, we observe an insignificant variation in $d_{10}^{*}$ evolution at all $We$ (see inset of Fig. 10a), it suggests that the the overall size distributions is prominently determined by the scale of $n$ maxima (see inset of Fig. 10b).
@@ IV Conclusion
In the present study, we have a closer look into the intermediate sub-scale processes that occur during the unsteady aerobreakup of a droplet in the range $We\approx 900-4000$ using a shock tube system. The subsequent multi-scale interfacial deformations, originating from various unstable wave mechanisms (KHI, RTI and RPI), forms a non-linear cascade. The undulations disintegrate through sub-secondary breakup processes, primarily conforming to a ligament or bag mode. The daughter droplet size distributions were estimated using a DFD technique, where the normalized distributions, at different $We$ and time periods, followed a common fit. This compound gamma distribution parameter aligns well with the coefficient determined from terminal ligaments. This corresponds to the limit associated with most extreme corrugations physically possible. This illustrates a self-similar behavior pertaining the size distributions as well as the evolution of topological features at different scales (initial droplet/undulation; global/local) with an effective $We$ bridging them.
@@ Acknowledgements.
S.J.R. acknowledges the support through the Prime Minister’s Research Fellowship (PMRF).

### v2

@@ III.3 Ligament-mediated droplet generation
As explained earlier, the sub-secondary breakup processes ends as the undulation disintegrates into daughter droplets. However, the termination entails a final ligament-mediated mechanism as illustrated in figure 4. The corrugated ligaments portrays an inertio-capillary system with liquid inertia and surface tension forces dominating the dynamics. The ligament can be considered as a collection of a protoblobs (see figure 7) where the evolution ensure two possibilities: either the blobs separate through RPI mechanism or they coalesce into a bigger blob. Recent studies [15, 27, 14] considered these effects and emphasized the role of corrugations on this delicate balance between breakup and coalescence. These random corrugations arise from the disturbances induced at extreme flow conditions and coupled interaction between the two phases. A correlation between the ligament shape and the final daughter droplet size distribution has been established in the form of a Gamma distribution [15] or possibly a compound gamma distribution [16]
@@
where $\Gamma$ is the gamma function, $\mathcal{K}$ is the modified Bessel function of the second kind and the coefficients $m$ and $n$ represent ligament size distribution and corrugations. The coefficient $n=\langle d_{c}\rangle^{2}/(\langle d_{c}^{2}\rangle-\langle d_{c}\rangle^{2})$ is deduced by decomposing a corrugated ligament into blobs matching local ligament diameter $d_{c}$ as illustrated in figure 7. The coefficient $m=\langle l\rangle^{2}/(\langle l^{2}\rangle-\langle l\rangle^{2})$ is deduced by taking the mean diameters $l=\langle d_{c}\rangle$ for various ligaments.
Considering a sample of ligaments spread across space and time with different imposed gas flow $We$, this corrugation factor was determined mostly to be in the range spanning $n=3-8$. The values close to $n=4$ were identified as the most frequently observed across various $We$ flows. This corresponds to extremely corrugated ligaments and $n=4$ is a practical limit on the coefficient considering various physical constraints [27]. Also a very low value of $m=3$ (corresponding to the fit as illustrated in figure 9a) represent a broad ligament size ($l$) distribution, as observed from the experiments as well (see figure 5). This illustrates the multi-scaled nature of the atomization processes generating the intermediate features at a range of scales.
The invariance with respect to $We$ number supports the saturation of these coefficients (especially $n$) to the theoretical limit. Although the sub-secondary breakup process involves the ligament production through various mechanisms, the ligaments collectively are postulated to exhibit extreme levels of corrugations, primarily due to the highly potent disruptive aerodynamic forces at extreme $We$. The ligaments are expected to scale down to smaller average diameters as $We$ is increased (as observed in the experiments in figure 5), but the shape factor is already saturated and depicts values close to $n=4$. To elucidate this hypothesis, we will consider the estimated size distribution in the subsequent sections.
@@ III.4 Droplet size distribution
The unsteady imposition arising from this open nozzle-blast wave induced flows [28, 18, 29], the dynamics are expected to be transient. However the droplet aerobreakup intrinsically is a transient phenomenon. The recurrent wave cascade continuously strips away the liquid from the parent droplet mass and generates droplets through sub-secondary breakup processes. Hence the droplets are produced with evolving characteristics over a period of time depending on the flow around residual parent droplet. To capture this effect, droplet size measurements were performed considering the time stamp of their appearance in the measurement zone ($t^{*}$). This zone was located $30mm$ downstream from the parent drop for a better visibility of daughter droplets[20].
@@
Figure 8: (a) Overall droplet size distribution PDFs at different $We$. (b) Temporal evolution of PDFs represented a contour plot, where the breakup event is sub divided into 5 equal periods. (c) Temporal evolution of Normalised PDFs (using average droplet diameter $\langle d\langle$), depicting self similar behaviour in time. Here $\tau_{b}$=centre of each time period/total breakup time.
@@
All the sampled droplets, when consolidated, gave the overall size distributions for different $We$ flow. However, the appearance of the first and the last droplet in the zone reflects the breakup time period $t_{b}^{*}$, which can be subdivided into equal sub-periods. The assessment of size distributions in these periods enables the determination of transient characteristics of the system. The time scale $t^{*}$, measured from the appearance of the first droplet, is then normalized using the breakup period $t_{b}^{*}$ such that $\tau_{b}=t^{*}/t_{b}^{*}$. The overall distributions as probability density functions (PDFs) are illustrated in figure 8a. The average droplet size decreases as $We$ is increased. The transient PDFs are represented as a contour plot depicting a transition to observing larger droplet sizes as time progresses in figure 8b. However, if we normalize the PDFs with corresponding average droplet size $d_{10}^{*}$, we observe the different time period to follows the same underlying distribution. The same effects are visible in all $We$ discussed in this study as illustrated in figure 8c.
@@
For more reliable statistics with a larger sample size in each period, the breakup time is split into 3 parts, and the corresponding normalized PDFs along with global distributions are presented in in figure 9a along with the compound gamma and log-normal distribution fits. They almost overlap, especially till $d/\langle d\rangle\approx 6$. The compound gamma distribution aligns with the earlier prediction $n=4$, deduced from the ligament-mediated mechanism. The deviation is observed consistently at the tails, which might be due to the effects like (i) The unsteady decaying airflow with lower effective $We$ at later stages results in larger droplets being produced from the residual parent droplet and (ii) coalescence of daughter droplets. To have a better understanding, we require unsteady destabilization models and better measurement capabilities.
@@
Since the normalized PDFs follow a common underlying distribution, the prediction of average diameter is ideally sufficient to determine actual PDFs. Sultanov and Yarin [30] assumed a chaotic isotropic deformation of the liquid mass to analytically deduce size distributions in explosive breakups, pertain high level of disruptive energy. We extend the same to present catastrophic aerobreakup. The imposed disruptive energy $E$ induces a deformation rate of $\dot{\gamma}$ to the droplet of initial radius $r_{0}$, such that $(1/2)\rho_{l}r_{0}^{2}(\dot{\gamma}r_{0})^{2}\sim E$. Assuming chaotic random dispersal [30], the same deformation rate is extended to individual daughter droplets with characteristic size $r_{f}$ (radius), such that the associated incremental surface energy scales as $(1/2)\rho_{l}r_{f}^{2}(\dot{\gamma}r_{0})^{2}\sim\sigma r_{f}^{2}$. Simplifying these equations where $E\sim\rho_{a}u_{a}^{2}r_{0}^{3}$ for aerobreakup, we get $(r_{f}/r_{0})\sim{We}^{-1/3}$. Suspecting $d_{10}\sim d_{f}$, we have
@@
This scaling relationship is validated in figure 9b. Although the observed breakup is not completely random, the hypothesis holds. Deviations from this are expected at lower $We$, where the breakup is evidently regular in geometry and prominently anisotropic. A similar scaling law for average droplet diameters was deduced by Kooij et al. [16] for flat-fan sprays, using the unstable wave-mediated sheet disintegration mechanism.
A temporal evolution of average droplet diameters $d_{10}^{*}$ are also presented in figure 10a, where global averages $d_{10}$ is used for normalization, depicting a similar trend for different $We$. The smaller sized droplets appear first (also see figure 8b) and the possible mechanisms include: (i) unsteady impulsive airflow with higher effective $We$ generating smaller droplets during early interaction (ii) size dependent migration where smaller droplets with lower inertia reach the measurement zone first. The number of daughter droplets $N$ can be also deduced extending earlier analysis
@@
We consider the number of droplets produced per unit measurement volume i.e. the number concentration $n=N/V_{meas}$ for unbiased estimates of droplet count [20, 31]. Temporal evolution of the normalized version $n/We$ is plotted in figure 10b illustrating the validity of aforementioned scales (represented by peaks). A similar trend is observed for different $We$, where a huge number of droplets are generated in the early stages. Culminating both the ideas (temporal evolution of $d_{10}^{*}$ and $n$), we deduce that the droplets with smaller diameters appear first in large quantities. Since, we observe an insignificant variation in temporal evolution of $d_{10}^{*}$ at all $We$ (see inset of figure 10a), it suggests that the the overall size distributions is predominantly determined by the scale associated with the maxima of $n$ (see inset of figure 10b).
@@ IV Conclusion
In the present study, we have a closer look into the intermediate sub-scale processes that occur during the unsteady aerobreakup of a droplet in the range $We\approx 900-4000$ using a shock tube system. The subsequent multi-scale interfacial deformations, originating from various unstable wave mechanisms (KHI, RTI and RPI), forms a non-linear cascade. The undulations disintegrate through sub-secondary breakup processes, primarily conforming to a ligament or bag mode. The daughter droplet size distributions were estimated using a DFD technique, where the normalized distributions, at different $We$ and time periods, followed a common fit. This compound gamma distribution parameter aligns well with the coefficient determined from terminal ligaments. This corresponds to the limit associated with most extreme corrugations physically possible. This illustrates a self-similar behavior pertaining the transient size distributions at different $We$, as well as the evolution of topological features at different scales (initial droplet/undulation; global/local) with an effective $We$ bridging them.
@@ Acknowledgements.
S.J.R. would like to thank the Prime Minister Research Fellowship (PMRF) for the financial support. S.B. would like to acknowledge the support from the Indian National Academy of Engineering (INAE) Chair professorship.